1.)Suppose I collect data on the average price of Yuengling (a type of beer) at every bar in Morgantown. I go to 6 bars and find an average of $3 with a sample standard deviation of $.60. If I wanted to test if the average cost of a Yuengling was equal to $3.50, what would my test statistic be? Round to two decimal places.

2.)The average length of the 13 songs on the Tame Impala album "Currents" is 236 seconds with a sample standard deviation of 110 seconds. I want to test if the average Currents song is longer than 200 seconds. What would my test statistic be? Round to two decimal places.

3.)Suppose I get a test statistic z = 2.34. If I am using a one tailed test, what is my p value? You should not have to round if you are using the table. If you are not, round to four decimal points.

4.)I am doing a two tailed z test with a z statistic of 1.72, what is my p value?

5.) I have a z statistic of 2.91 from a one tailed test. What is my p value?
Answer #1

Ans:

1)

Test statistic:

t=(3-3.50)/(0.60/sqrt(6))

t=-2.04

2)

Test statistic:

t=(236-200)/(110/SQRT(13)

t=1.18

3)One tailed test

p-value=P(z>2.34)=0.0096

4)Two tailed test

p-value=2*P(z>1.72)=0.0854

5)One tailed test

p-value=P(z>2.91)=0.0018